# Measuring Self-Rating Bias in LLM-Generated Survey Data: A Semantic Similarity Framework for Independent Scale Mapping

### v1

@@ Abstract
Synthetic survey data generated by large language models (LLMs) suffers from a fundamental circularity: the same model family that generates text responses also maps them to numerical scales. We calibrate and validate Semantic Similarity Rating (SSR; maier2024ssr), which decouples generation from scale mapping via embedding-based cosine similarity against predefined anchor statements. Configuration experiments ($N=17$ pilot, $N=69$ cross-validation across 8 domains) show that naturalistic behavioral anchors outperform formal jargon by 29 percentage points (pp), and that SSR achieves 65–67% exact match and 91% within $\pm$1; a cross-model test with OpenAI text-embedding-3-small reaches 77% exact, confirming cross-provider generalization. Direct LLM baselines (Claude 87%, GPT-4o 83%) establish that SSR’s contribution is methodological independence, not accuracy superiority. A control condition removing question text from the LLM prompt actually improves LLM accuracy, ruling out information asymmetry as the explanation for SSR’s lower accuracy. A pre-registered circularity experiment ($N=345$) reveals 4-fold compressed error variance in LLM rating ($\sigma^{2}=0.21$ vs. $0.87$ for SSR) and systematic directional bias. A cross-model control (GPT-4o rating Claude-generated text) shows nearly identical compression (within/cross ratio =0.93=0.93), indicating variance compression is a general LLM property rather than a within-model artifact. The calibration dataset, anchor library, and source code are publicly available (see Data Availability).
keywords:
synthetic survey data, semantic similarity, text embeddings, Likert scales, language models, survey methodology
@@ 1.1 The Promise and Peril of Synthetic Survey Data
Survey research is the backbone of the behavioral and social sciences, yet it faces mounting challenges: declining response rates (groves2009survey), rising costs, and increasing difficulty reaching representative samples. Against this backdrop, large language models (LLMs) have emerged as a promising tool for generating synthetic survey responses, simulated data that mimics human respondents’ answers to structured questionnaire items (argyle2023out; horton2023llm).
If LLMs can produce realistic survey data that mirrors the distributional properties of real human samples, researchers could use synthetic data for instrument pretesting, power analysis, hypothesis generation, and, most ambitiously, as a complement to or partial substitute for expensive human data collection (park2023generative). Several recent studies have demonstrated that LLM-generated responses can reproduce demographic patterns in political attitudes (argyle2023out), replicate findings from behavioral economics experiments (horton2023llm; aher2023using), and generate plausible responses to standard survey instruments (bisbee2024synthetic).
@@ 1.2 The Circularity Problem
This is not independent measurement. It is the model evaluating its own output, analogous to asking a student to grade their own exam. The resulting data conflates two distinct measurement operations that psychometric theory requires to be independent: the generation of a response and the scoring of that response on a measurement instrument (cronbach1955construct).
The consequences of this circularity are subtle but consequential. Reliability metrics computed on circularly-validated data will be inflated, because the model’s internal consistency reflects its own biases rather than the construct being measured. Validity claims based on agreement between LLM-generated text and LLM-assigned ratings are fundamentally non-transferable, because they measure only the coherence of the model’s own outputs rather than alignment with human judgment (maier2024ssr).
@@ 1.3 Our Contribution
We systematically calibrate and validate Semantic Similarity Rating (SSR; (maier2024ssr)), a framework that addresses circularity by decoupling text generation from scale mapping, through four key contributions:
(1) we implement and validate SSR using Voyage AI voyage-3.5-lite (voyageai2024embeddings) as an architecturally independent measurement channel, confirming cross-provider generalization with OpenAI text-embedding-3-small (77% vs. 67% exact match); (2) we demonstrate that naturalistic behavioral anchors outperform formal survey jargon by 29 percentage points (pp); (3) we provide systematic calibration through factorial experiments identifying the optimal configuration; and (4) a pre-registered circularity experiment ($N=345$) with cross-model control reveals 4-fold variance compression in LLM-based rating, established as a general LLM property (within/cross ratio =0.93=0.93) rather than a within-model artifact.
@@ 2.1 LLMs as Survey Respondents
The use of LLMs to simulate survey respondents has rapidly expanded since the seminal work of argyle2023out, who demonstrated that GPT-3 could produce response distributions conditioned on demographic variables that closely matched those of real American subpopulations across political, economic, and social attitude items. Subsequent work has extended this approach to economic decision-making (horton2023llm), social behavior simulation (park2023generative; aher2023using), and political survey replication (bisbee2024synthetic).
@@
However, critical evaluations have revealed important limitations. santurkar2023whose showed that LLM-generated opinion distributions are systematically biased toward certain demographic groups, reflecting the composition and biases of training data rather than any genuine model of human attitudes. bisbee2024synthetic demonstrated that synthetic survey replacements can produce misleading results on novel political questions, particularly when the question format or content departs from the model’s training distribution.
@@
A critical gap in this literature is the relative inattention to the measurement pipeline, specifically how text responses are mapped to numeric scale values. Most studies either instruct the LLM to produce numeric output directly, parse structured responses, or use the same (or similar) model to rate its own output. The SSR framework proposed by maier2024ssr represents the closest precursor to our work. They demonstrated that embedding-based semantic similarity can map LLM-generated text to Likert distributions that closely replicate real human survey data, testing on 57 personal care product surveys with 9,300 human responses as ground truth from a leading consumer goods corporation. Their validation is distributional: they show that synthetic response distributions match human distributions at the population level (Kolmogorov–Smirnov similarity >0.85>0.85, achieving 90% of human test–retest reliability). This is a powerful result that establishes SSR’s viability for applied market research, but it addresses a different question than individual-level accuracy: distributional similarity can hold even when individual predictions contain errors, because overestimates and underestimates cancel at the aggregate level.
@@
Our work complements theirs by addressing the engineering and methodological questions that distributional validation leaves open. First, we provide systematic calibration through factorial experiments that identify why SSR works—discovering that anchor statement quality (+29 pp) and asymmetric embedding (+6 pp) are the dominant factors—rather than demonstrating that it works at the distributional level. Second, we frame the contribution explicitly in terms of the circularity problem: the measurement independence that embedding-based mapping provides relative to LLM self-rating. Notably, maier2024ssr’s comparison of follow-up Likert rating (where the same model generates and rates text) against SSR shows inferior distributional similarity for the self-rating approach—a result consistent with our finding that self-rating circularity introduces measurable bias—but they do not identify circularity as the mechanism driving this difference. Third, we generalize beyond their purchase intent focus to 8 semantic domains, revealing substantial domain-specific variation (33–90% exact match) that motivates targeted anchor refinement. Finally, we compare SSR against a direct LLM baseline with controlled methodology (temperature =0=0, two prompt variants), establishing the accuracy–independence tradeoff that practitioners must navigate.
@@ 2.2 Text Embeddings and Semantic Similarity
Text embeddings, dense vector representations of natural language that encode semantic content in continuous space, provide the mathematical foundation for SSR. Modern embedding models produce vectors where cosine similarity correlates with semantic relatedness (reimers2019sentence).
The quality of these representations has improved substantially with advances in contrastive learning. gao2021simcse introduced SimCSE, demonstrating that simple contrastive objectives could substantially improve sentence embedding quality. The Massive Text Embedding Benchmark (muennighoff2023mteb) established standardized evaluation across retrieval, classification, and semantic similarity tasks, enabling systematic comparison of embedding models.
@@
Modern embedding models also support asymmetric embedding, where query and document texts are embedded with different objectives (wang2022text). This is standard practice in information retrieval, where queries and documents have different distributional properties. We exploit this capability by embedding response texts as queries and anchor statements as documents, treating scale mapping as a retrieval problem.
@@ 2.3 Anchor-Based Measurement in Psychology
The design of anchor statements for Likert scales has a long history in psychometrics (likert1932technique). krosnick2010question established that the specific language of scale endpoints and anchors significantly affects response behavior, with behavioral descriptions producing more consistent and valid responses than abstract labels.
tourangeau2000psychology proposed a cognitive model of survey response in which respondents (1) comprehend the question, (2) retrieve relevant information, (3) form a judgment, and (4) map that judgment to the response scale. The fourth step, response mapping, is where anchor language exerts its influence. Clear, concrete anchors reduce mapping ambiguity and improve measurement precision.
@@ 3.3.1 Embedding Model
We use Voyage AI’s voyage-3.5-lite model as the primary embedding provider, which produces 1024-dimensional vectors trained for semantic similarity and information retrieval tasks. We selected this model for three reasons: (1) it ranks competitively on the Massive Text Embedding Benchmark (MTEB; muennighoff2023mteb) across retrieval and semantic similarity tasks; (2) it natively supports asymmetric embedding via an input_type parameter that optimizes vectors differently for document and query roles, which we exploit for scale mapping; and (3) it offers favorable cost characteristics (\$0.02 per million tokens) suitable for production deployment. To verify that SSR’s effectiveness is not specific to Voyage, we also evaluate with OpenAI’s text-embedding-3-small (1536 dimensions), which does not support asymmetric embedding, providing a fair symmetric-to-symmetric comparison (Section 4.8).
@@ 3.4.1 The Compression Problem
Raw cosine similarities between text embeddings cluster in a narrow range—typically spanning only 0.05–0.10 units (e.g., $[0.75,0.83]$ or $[0.81,0.90]$ depending on the specific texts and embedding model). This compression phenomenon, a known property of high-dimensional embedding spaces, means that a naïve softmax over raw similarities produces near-uniform distributions that lack discriminative power. Figure 2 illustrates this problem on a representative test case.
@@ 3.6 Model and API Details
Table 1 lists the specific models and API parameters used throughout this study. Text generation and LLM rating used Claude Haiku 4.5 (anthropic2024claude3); cross-model validation used GPT-4o (openai2023gpt4); embeddings used Voyage 3.5-lite (voyageai2024embeddings) and OpenAI text-embedding-3-small. All API calls were made between January and February 2026.
@@ 4.4.3 Findings
  • H2 (context) is harmful: Adding question text to anchor statements reduces accuracy by 17 percentage points (82% $\to$ 65%). The question text introduces shared semantic content that makes all anchors more similar to each other, reducing the discriminability between scale points. Alternative contextualization strategies—such as embedding question and response jointly while keeping anchors context-free, or using learned question-conditioned anchor weighting—were not explored and may yield different results.
@@ 4.5 Final Configuration (Pilot Result)
  • Temperature: $\tau=0.2$ (pilot default; subsequent cross-validation on 8 domains identified $\tau=0.15$ as optimal—see Section 4.6. We report results at both values throughout for transparency)
@@ 4.7 Baseline Comparison: Direct LLM Rating
Both LLM variants achieve identical global accuracy (87% exact, 100% within $\pm$1, MAE =0.13=0.13), though they differ in which cases they miss: the as-deployed variant errs on 9 cases while the domain-aware variant errs on a different set of 9 cases. All LLM errors across both variants are off-by-one, and 8 of 9 errors in each variant are underestimates (predicted one point below expected), indicating a slight negative bias for ambiguous cases. The convergence in global performance despite different error patterns suggests that scale label specificity does not substantially affect LLM rating accuracy—the model’s sentiment analysis capability is robust to label framing.
4.7.1 Information Asymmetry — Control Condition
An important methodological difference between the two approaches is that the LLM baseline receives the full question text as part of its prompt, while SSR does not incorporate question text during similarity computation. To quantify whether this information asymmetry explains the accuracy gap, we ran a control condition: both models (Claude Haiku 4.5 and GPT-4o) rating the same 69 test cases with the question text removed from the prompt, providing only the response text and scale anchors—the same information available to SSR. Table 14 reports the results.
@@ 4.8 Cross-Model Generalization
The preceding experiments used a single embedding model (Voyage 3.5-lite) and a single LLM family (Anthropic Claude) for baseline comparison. To assess whether our findings generalize across model providers, we evaluated both the SSR framework and the LLM baseline using alternative models: OpenAI text-embedding-3-small (1536 dimensions) for embedding-based mapping and GPT-4o (temperature =0=0) for direct LLM rating. This directly addresses the concern that SSR’s accuracy or the baseline comparison could be artifacts of a specific model’s characteristics, and follows the multi-model approach of maier2024ssr, who validated SSR with GPT-4o, Gemini-2f, and OpenAI embeddings.
@@ 4.8.1 Embedding Model Comparison
OpenAI’s embedding model substantially outperforms Voyage in symmetric mode (+12 pp exact match, +4 pp within $\pm$1, 0.18-0.18 MAE). Figure 7 visualizes the per-domain comparison: OpenAI achieves 100% exact match on ease—the weakest Voyage domain—and 88% on likelihood, agreement, value, and purchase intent.
@@
Critically, OpenAI achieves 100% within $\pm$1 across all temperatures tested ($\tau\in\{0.05,0.10,0.15,0.20,0.25\}$), compared to Voyage symmetric’s 88–97%. This indicates that the SSR framework’s ability to identify the correct scale region generalizes robustly across embedding providers.

### v2

@@ Abstract
Synthetic survey data generated by large language models (LLMs) suffers from a fundamental circularity: the same model family that generates text responses also maps them to numerical scales. We calibrate and validate Semantic Similarity Rating (SSR; Maier et al., 2024), which decouples generation from scale mapping via embedding-based cosine similarity against predefined anchor statements. Configuration experiments ($N=17$ pilot, $N=69$ cross-validation across 8 domains) show that naturalistic behavioral anchors outperform formal jargon by 29 percentage points (pp), and that SSR achieves 65–67% exact match and 91% within $\pm$1; a cross-model test with OpenAI text-embedding-3-small reaches 77% exact, confirming cross-provider generalization. Direct LLM baselines (Claude 87%, GPT-4o 83%) establish that SSR’s contribution is methodological independence, not accuracy superiority. A control condition removing question text from the LLM prompt actually improves LLM accuracy, ruling out information asymmetry as the explanation for SSR’s lower accuracy. A pre-registered circularity experiment ($N=345$) reveals 4-fold compressed error variance in LLM rating ($\sigma^{2}=0.21$ vs. $0.87$ for SSR) and systematic directional bias. A cross-model control (GPT-4o rating Claude-generated text) shows nearly identical compression (within/cross ratio =0.93=0.93), indicating variance compression is a general LLM property rather than a within-model artifact. The calibration dataset, anchor library, and source code are publicly available (see Data Availability).
Keywords: synthetic survey data, semantic similarity, text embeddings, Likert scales, language models, survey methodology
@@ 1.1 The Promise and Peril of Synthetic Survey Data
Survey research is the backbone of the behavioral and social sciences, yet it faces mounting challenges: declining response rates (Groves et al.,, 2009), rising costs, and increasing difficulty reaching representative samples. Against this backdrop, large language models (LLMs) have emerged as a promising tool for generating synthetic survey responses, simulated data that mimics human respondents’ answers to structured questionnaire items (Argyle et al.,, 2023; Horton,, 2023).
If LLMs can produce realistic survey data that mirrors the distributional properties of real human samples, researchers could use synthetic data for instrument pretesting, power analysis, hypothesis generation, and, most ambitiously, as a complement to or partial substitute for expensive human data collection (Park et al.,, 2023). Several recent studies have demonstrated that LLM-generated responses can reproduce demographic patterns in political attitudes (Argyle et al.,, 2023), replicate findings from behavioral economics experiments (Horton,, 2023; Aher et al.,, 2023), and generate plausible responses to standard survey instruments (Bisbee et al.,, 2024).
@@ 1.2 The Circularity Problem
This is not independent measurement. It is the model evaluating its own output, analogous to asking a student to grade their own exam. The resulting data conflates two distinct measurement operations that psychometric theory requires to be independent: the generation of a response and the scoring of that response on a measurement instrument (Cronbach and Meehl,, 1955).
The consequences of this circularity are subtle but consequential. Reliability metrics computed on circularly-validated data will be inflated, because the model’s internal consistency reflects its own biases rather than the construct being measured. Validity claims based on agreement between LLM-generated text and LLM-assigned ratings are fundamentally non-transferable, because they measure only the coherence of the model’s own outputs rather than alignment with human judgment (Maier et al.,, 2024).
@@ 1.3 Our Contribution
We systematically calibrate and validate Semantic Similarity Rating (SSR; (Maier et al.,, 2024)), a framework that addresses circularity by decoupling text generation from scale mapping, through four key contributions:
(1) we implement and validate SSR using Voyage AI voyage-3.5-lite (Voyage AI,, 2024) as an architecturally independent measurement channel, confirming cross-provider generalization with OpenAI text-embedding-3-small (77% vs. 67% exact match); (2) we demonstrate that naturalistic behavioral anchors outperform formal survey jargon by 29 percentage points (pp); (3) we provide systematic calibration through factorial experiments identifying the optimal configuration; and (4) a pre-registered circularity experiment ($N=345$) with cross-model control reveals 4-fold variance compression in LLM-based rating, established as a general LLM property (within/cross ratio =0.93=0.93) rather than a within-model artifact.
@@ 2.1 LLMs as Survey Respondents
The use of LLMs to simulate survey respondents has rapidly expanded since the seminal work of Argyle et al., (2023), who demonstrated that GPT-3 could produce response distributions conditioned on demographic variables that closely matched those of real American subpopulations across political, economic, and social attitude items. Subsequent work has extended this approach to economic decision-making (Horton,, 2023), social behavior simulation (Park et al.,, 2023; Aher et al.,, 2023), and political survey replication (Bisbee et al.,, 2024).
@@
However, critical evaluations have revealed important limitations. Santurkar et al., (2023) showed that LLM-generated opinion distributions are systematically biased toward certain demographic groups, reflecting the composition and biases of training data rather than any genuine model of human attitudes. Bisbee et al., (2024) demonstrated that synthetic survey replacements can produce misleading results on novel political questions, particularly when the question format or content departs from the model’s training distribution.
@@
A critical gap in this literature is the relative inattention to the measurement pipeline, specifically how text responses are mapped to numeric scale values. Most studies either instruct the LLM to produce numeric output directly, parse structured responses, or use the same (or similar) model to rate its own output. The SSR framework proposed by Maier et al., (2024) represents the closest precursor to our work. They demonstrated that embedding-based semantic similarity can map LLM-generated text to Likert distributions that closely replicate real human survey data, testing on 57 personal care product surveys with 9,300 human responses as ground truth from a leading consumer goods corporation. Their validation is distributional: they show that synthetic response distributions match human distributions at the population level (Kolmogorov–Smirnov similarity >0.85>0.85, achieving 90% of human test–retest reliability). This is a powerful result that establishes SSR’s viability for applied market research, but it addresses a different question than individual-level accuracy: distributional similarity can hold even when individual predictions contain errors, because overestimates and underestimates cancel at the aggregate level.
@@
Our work complements theirs by addressing the engineering and methodological questions that distributional validation leaves open. First, we provide systematic calibration through factorial experiments that identify why SSR works, discovering that anchor statement quality (+29 pp) and asymmetric embedding (+6 pp) are the dominant factors, rather than demonstrating that it works at the distributional level. Second, we frame the contribution explicitly in terms of the circularity problem: the measurement independence that embedding-based mapping provides relative to LLM self-rating. Notably, Maier et al.,’s comparison of follow-up Likert rating (where the same model generates and rates text) against SSR shows inferior distributional similarity for the self-rating approach, a result consistent with our finding that self-rating circularity introduces measurable bias, but they do not identify circularity as the mechanism driving this difference. Third, we generalize beyond their purchase intent focus to 8 semantic domains, revealing substantial domain-specific variation (33–90% exact match) that motivates targeted anchor refinement. Finally, we compare SSR against a direct LLM baseline with controlled methodology (temperature =0=0, two prompt variants), establishing the accuracy–independence tradeoff that practitioners must navigate.
@@ 2.2 Text Embeddings and Semantic Similarity
Text embeddings, dense vector representations of natural language that encode semantic content in continuous space, provide the mathematical foundation for SSR. Modern embedding models produce vectors where cosine similarity correlates with semantic relatedness (Reimers and Gurevych,, 2019).
The quality of these representations has improved substantially with advances in contrastive learning. Gao et al., (2021) introduced SimCSE, demonstrating that simple contrastive objectives could substantially improve sentence embedding quality. The Massive Text Embedding Benchmark (Muennighoff et al.,, 2023) established standardized evaluation across retrieval, classification, and semantic similarity tasks, enabling systematic comparison of embedding models.
@@
Modern embedding models also support asymmetric embedding, where query and document texts are embedded with different objectives (Wang et al.,, 2022). This is standard practice in information retrieval, where queries and documents have different distributional properties. We exploit this capability by embedding response texts as queries and anchor statements as documents, treating scale mapping as a retrieval problem.
@@ 2.3 Anchor-Based Measurement in Psychology
The design of anchor statements for Likert scales has a long history in psychometrics (Likert,, 1932). Krosnick and Presser, (2010) established that the specific language of scale endpoints and anchors significantly affects response behavior, with behavioral descriptions producing more consistent and valid responses than abstract labels.
Tourangeau et al., (2000) proposed a cognitive model of survey response in which respondents (1) comprehend the question, (2) retrieve relevant information, (3) form a judgment, and (4) map that judgment to the response scale. The fourth step, response mapping, is where anchor language exerts its influence. Clear, concrete anchors reduce mapping ambiguity and improve measurement precision.
@@ 3.3.1 Embedding Model
We use Voyage AI’s voyage-3.5-lite model as the primary embedding provider, which produces 1024-dimensional vectors trained for semantic similarity and information retrieval tasks. We selected this model for three reasons: (1) it ranks competitively on the Massive Text Embedding Benchmark (MTEB; Muennighoff et al.,, 2023) across retrieval and semantic similarity tasks; (2) it natively supports asymmetric embedding via an input_type parameter that optimizes vectors differently for document and query roles, which we exploit for scale mapping; and (3) it offers favorable cost characteristics (\$0.02 per million tokens) suitable for production deployment. To verify that SSR’s effectiveness is not specific to Voyage, we also evaluate with OpenAI’s text-embedding-3-small (1536 dimensions), which does not support asymmetric embedding, providing a fair symmetric-to-symmetric comparison (Section 4.8).
@@ 3.4.1 The Compression Problem
Raw cosine similarities between text embeddings cluster in a narrow range, typically spanning only 0.05–0.10 units (e.g., $[0.75,0.83]$ or $[0.81,0.90]$ depending on the specific texts and embedding model). This compression phenomenon, a known property of high-dimensional embedding spaces, means that a naïve softmax over raw similarities produces near-uniform distributions that lack discriminative power. Figure 2 illustrates this problem on a representative test case.
@@ 3.6 Model and API Details
Table 1 lists the specific models and API parameters used throughout this study. Text generation and LLM rating used Claude Haiku 4.5 (Anthropic,, 2024); cross-model validation used GPT-4o (OpenAI,, 2023); embeddings used Voyage 3.5-lite (Voyage AI,, 2024) and OpenAI text-embedding-3-small. All API calls were made between January and February 2026.
@@ 4.4.3 Findings
  • H2 (context) is harmful: Adding question text to anchor statements reduces accuracy by 17 percentage points (82% $\to$ 65%). The question text introduces shared semantic content that makes all anchors more similar to each other, reducing the discriminability between scale points. Alternative contextualization strategies (such as embedding question and response jointly while keeping anchors context-free, or using learned question-conditioned anchor weighting) were not explored and may yield different results.
@@ 4.5 Final Configuration (Pilot Result)
  • Temperature: $\tau=0.2$ (pilot default; subsequent cross-validation on 8 domains identified $\tau=0.15$ as optimal; see Section 4.6. We report results at both values throughout for transparency)
@@ 4.7 Baseline Comparison: Direct LLM Rating
Both LLM variants achieve identical global accuracy (87% exact, 100% within $\pm$1, MAE =0.13=0.13), though they differ in which cases they miss: the as-deployed variant errs on 9 cases while the domain-aware variant errs on a different set of 9 cases. All LLM errors across both variants are off-by-one, and 8 of 9 errors in each variant are underestimates (predicted one point below expected), indicating a slight negative bias for ambiguous cases. The convergence in global performance despite different error patterns suggests that scale label specificity does not substantially affect LLM rating accuracy; the model’s sentiment analysis capability is robust to label framing.
4.7.1 Information Asymmetry: Control Condition
An important methodological difference between the two approaches is that the LLM baseline receives the full question text as part of its prompt, while SSR does not incorporate question text during similarity computation. To quantify whether this information asymmetry explains the accuracy gap, we ran a control condition: both models (Claude Haiku 4.5 and GPT-4o) rating the same 69 test cases with the question text removed from the prompt, providing only the response text and scale anchors (the same information available to SSR). Table 14 reports the results.
@@ 4.8 Cross-Model Generalization
The preceding experiments used a single embedding model (Voyage 3.5-lite) and a single LLM family (Anthropic Claude) for baseline comparison. To assess whether our findings generalize across model providers, we evaluated both the SSR framework and the LLM baseline using alternative models: OpenAI text-embedding-3-small (1536 dimensions) for embedding-based mapping and GPT-4o (temperature =0=0) for direct LLM rating. This directly addresses the concern that SSR’s accuracy or the baseline comparison could be artifacts of a specific model’s characteristics, and follows the multi-model approach of Maier et al., (2024), who validated SSR with GPT-4o, Gemini-2f, and OpenAI embeddings.
@@ 4.8.1 Embedding Model Comparison
OpenAI’s embedding model substantially outperforms Voyage in symmetric mode (+12 pp exact match, +4 pp within $\pm$1, 0.18-0.18 MAE). Figure 7 visualizes the per-domain comparison: OpenAI achieves 100% exact match on ease (the weakest Voyage domain) and 88% on likelihood, agreement, value, and purchase intent.
@@
Critically, OpenAI achieves 100% within $\pm$1 across all five temperatures tested ($\tau$ from 0.05 to 0.25), compared to Voyage symmetric’s 88–97%. This indicates that the SSR framework’s ability to identify the correct scale region generalizes robustly across embedding providers.